# Smriti Singh*1, Shuvam Keshari*1, Vinija Jain2, Aman Chadha2,3

### v2

@@ 2.1 Dataset Design
The design of this dataset is ethically motivated by the question, "Does socioeconomic status (SES) contribute to the ability to empathize with socioeconomically underprivileged people?" Studies from psychology show that SES may be a factor in the ability to experience empathy Jolliffe and Farrington (2006); Love (2020). Further, the consensus from social media seems to be that the people on the higher end of the socioeconomic spectrum remain unanswered and don’t understand the challenges that the people on the opposite end of the spectrum face. However, the answer to this question, and the effect it may have remains unanswered as far as LLMs are concerned.
@@
Figure 3: Examples of topic classification of SilverSpoon.
@@ 2.3 Data Analysis
Figure 5: SilverSpoon data distribution. Refer to section 2.3 for information about each topic.
@@ 4.1 Quantitative Analysis
We report per class accuracy, Cohen’s Kappa, Jaccard Similarity, Dice Similarity, and Matthew’s Correlation Coefficients against the upper-end labels and lower-end labels respectively. We do not use F1, precision, and recall in our evaluation since these metrics are calculated against a true-positive class, and our goal is to understand the similarity between model responses and each category of labels. Table 1 presents these results, To convert the model responses to a binary scale, we employ automatic evaluation, asking GPT4 to assess various responses and assign them to Class 1: Supporting the socioeconomically underprivileged, or Class 0: Otherwise. The prompt for this can be found in the appendix E.6.
@@
The range of values for accuracy, JSC, and DSC is 0 to 1 whereas the range for CKS and MCC is -1 to 1. For the latter two metrics, 0 indicates random chance agreement. We find that Alpaca has a negative CKC and MCC implying that its correlation with the lower-end labels is less than random, hence it agrees more with the upper class labels. More information can be found in Appendix D
Figure 8: A bird’s eye view depicting the variation of all model responses favoring the socioeconomically underprivileged. Here, L.E. stands for lower-end labels and U.E. stands for upper-end labels.
@@ 4.2 Qualitative Analysis
To perform a high-caliber qualitative analysis, we evaluate each model while considering the topic distribution in Figure 5. We acknowledge that these topics may vary depending upon various models employed to find the topics, and this analysis is meant to act as a baseline that can augment future research endeavors that utilize this dataset.
@@ 4.2.1 GPT
Figure 9: Judge GPT-4? A depiction of how GPT4 almost always only seems to be concerned with the legal aspect of any scenario, including domestic violence. In the figure, L.E stands for lower-end labels and U.E stands for upper-end labels.
@@
Figure 10: Detox is wrong but stealing is okay? Examples of unusual responses by GPT-3.5. In the figure, L.E stands for lower-end labels and U.E stands for upper-end labels.
@@ 4.2.2 Llama-2-Chat
Figure 11: Both are wrong? Examples of unusual responses by Llama-13B. Here, L.E stands for lower-end labels and U.E stands for upper-end labels.
We find that Llama-2-Chat-13B says "both are wrong" more often than other models. (For quantitative evaluation, we discard these responses.) Llama-2-Chat-13B has a very similar response distribution to GPT-3.5, a fact that is reflected in Figure 8. The most frequently occurring explanation this model offers for siding with the underprivileged is along the lines of "Respecting others’ boundaries" is important. In this manner, Llama-2-Chat displays more sympathy than GPT-4 (but marginally lesser than GPT-3.5) toward the socioeconomically underprivileged. However, it is interesting to see how Llama2 reacts to different situations, and a few examples have been demonstrated in Figure 11. Noticeably, Refer to Figure 19 for the topic-wise analysis of model responses.
@@ 4.2.3 Alpaca
Figure 13: Everyone is wrong? Some examples that highlight Alpaca’s illogical responses. (L.E. stands for lower-end labels and U.E. stands for upper-end labels.)
@@ 4.2.4 Zephyr
Figure 14: A depiction of Zephyr’s ability to provide comprehensive responses. (L.E. stands for lower-end labels and U.E. stands for upper-end labels.)
@@
Figure 15: You are correct? An illustration of Incoherent responses by Yi-6B. (L.E. stands for lower-end labels and U.E stands for upper-end labels.)
@@
Figure 16: Poor people have no rights? Examples of unusual responses by Yi-6B. L.E. stands for lower-end labels and U.E. stands for upper-end labels.
@@ 6 Conclusion
In this paper, we present SilverSpoon, the first dataset (to the best of our knowledge) with dual label annotation that is designed to help investigate the presence of socioeconomic bias in SOTA LLMs, and their ability to empathize with socioeconomically underprivileged people in difficult situations. Further, we present a comprehensive analysis of SOTA LLMs and the socioeconomic bias they exhibit. We find that most LLMs are unable to exhibit any empathy for socioeconomically underprivileged people in difficult conditions. We also observe that Alpaca displays disagreement with the lower-end labels. This research is meant to act as a first step toward investigating socioeconomic bias in language models. Specifically, future work in this area could focus on expanding this dataset to have different types of questions, more annotators, or even questions in different cultural/linguistic contexts. We hope that this dataset paves the way for further research in this domain.
@@ Appendix C Model responses classified by topics
Figure 20: Llama-2-Chat-7B mostly agrees with GPT-4: Responses illustrated by topic. Refer to Figure 9 for comparison.

### v3

@@ 2.1 Dataset Design
The design of this dataset is ethically motivated by the question, "Does socioeconomic status (SES) contribute to the ability to empathize with socioeconomically underprivileged people?". Many studies from psychology and sociology show that SES may be a factor in the ability to experience empathy Liu et al. (2023); Pishghadam et al. (2022); Jolliffe and Farrington (2006); Li and Siu (2021); Love (2020). Further, the consensus from social media seems to be that the people on the higher end of the socioeconomic spectrum don’t understand the challenges that the people on the opposite end of the spectrum face  [^2] Reddit Thread 1, Reddit Thread 2, Reddit Thread 3. However, the answer to this question, and the effect it may have remains unanswered as far as LLMs are concerned.
@@
Figure 3: Examples of topic classification of our SilverSpoon dataset.
@@ 2.3 Data Analysis
Figure 5: SilverSpoon data distribution. Please refer to Section 2.3 for information about each topic.
@@ 4.1 Quantitative Analysis
We report per class accuracy, Cohen’s Kappa, Jaccard Similarity, Dice Similarity, and Matthew’s Correlation Coefficients against the upper-end labels and lower-end labels respectively. We do not use F1, precision, and recall in our evaluation since these metrics are calculated against a true-positive class, and our goal is to understand the similarity between model responses and each category of labels. We present multiple metrics in order to illustrate a complete picture of our quantitative analysis. Table 1 presents these results, To convert the model responses to a binary scale, we employ automatic evaluation, asking GPT4 to assess various responses and assign them to Class 1: Supporting the socioeconomically underprivileged, or Class 0: Otherwise. The prompt for this can be found in the appendix E.6.
@@
The range of values for accuracy, JSC, and DSC is 0 to 1 whereas the range for CKS and MCC is -1 to 1. For the latter two metrics, 0 indicates random chance agreement. We find that Alpaca has a negative CKC and MCC implying that its correlation with the lower-end labels is less than random, hence it agrees more with the upper class labels. AppendixD offers a detailed discourse on the metrics of our choice and what they signify.
Figure 8: A bird’s eye view depicting the variation of all model responses favoring the socioeconomically underprivileged.
@@ 4.2 Qualitative Analysis
To perform a high-caliber qualitative analysis, we evaluate each model while considering the topic distribution in Figure 5. We acknowledge that these topics may vary depending upon various models employed to find the topics, and this analysis is meant to act as a baseline that can augment future research endeavors that utilize this dataset. In Tables 9-16, L.E stands for "lower-end" labels and U.E stands for "higher-end" labels.
@@ 4.2.1 GPT
Figure 9: Judge GPT-4? A depiction of how GPT4 almost always only seems to be concerned with the legal aspect of any scenario, including domestic violence.
@@
Figure 10: Detox is wrong but stealing is OK? Examples of unusual responses by GPT-3.5.
@@ 4.2.2 Llama-2-Chat
Figure 11: Both are wrong? Examples of unusual responses by Llama-13B.
We find that Llama-2-Chat-13B says "both are wrong" more often than other models. (For quantitative evaluation, we discard these responses.) Llama-2-Chat-13B has a very similar response distribution to GPT-3.5, a fact that is reflected in Figure 8. The most frequently occurring explanation this model offers for siding with the underprivileged is along the lines of "Respecting others’ boundaries" is important. In this manner, Llama-2-Chat displays more sympathy than GPT-4 (but marginally lesser than GPT-3.5) toward the socioeconomically underprivileged. However, it is interesting to see how Llama2 reacts to different situations, and a few examples have been demonstrated in Figure 11. Figure 19 shows a topic-wise analysis of model responses.
@@ 4.2.3 Alpaca
Figure 13: Everyone is wrong? Some examples that highlight Alpaca’s illogical responses.
@@ 4.2.4 Zephyr
Figure 14: A depiction of Zephyr’s ability to provide comprehensive responses.
@@
Figure 15: You are correct? An illustration of Incoherent responses by Yi-6B.
@@
Figure 16: Poor people have no rights? Examples of unusual responses by Yi-6B.
@@ 6 Conclusion
In this paper, we present SilverSpoon, the first dataset with dual label annotation that is designed to help investigate the presence of socioeconomic bias in SoTA LLMs, and their ability to empathize with socioeconomically underprivileged people in difficult situations. Further, we present a comprehensive analysis of SOTA LLMs and the socioeconomic bias they exhibit. We find that most LLMs are unable to exhibit any empathy for socioeconomically underprivileged people in difficult conditions. We also observe that Alpaca displays disagreement with the lower-end labels. This research is meant to act as a first step toward investigating socioeconomic bias in language models. Specifically, future work in this area could focus on expanding this dataset to have different types of questions, more annotators, or even questions in different cultural/linguistic contexts. We hope that this dataset paves the way for further research in this domain.
@@ Appendix C Model responses classified by topics
Figure 20: Llama-2-Chat-7B mostly agrees with GPT-4: Responses illustrated by topic. Refer to Figure LABEL:fig:GPT4_answers for comparison.